2y^-2 -3y^-1 -2 =0

Answer by troymuga(3)

2/(y^2) - 3/y = 2
ALL * y
2/y - 3 = 2y
Switch'em around
2/y - 2y = 3
ALL *y
2 - 2y^2 = 3y
ARRANGE THEM TO ONE SIDE
2y^2 +3y - 2
FACTOR the POLYNOMIAL
(2y - 1)(y + 2) = 0
SOLVE for y
y = 1/2, y = -2